Cherry Blossom Products Inc. produces and sells yoga-training products: how-to DVDs and a basic equipment set (blocks, strap, and small pillows). Last year, Cherry Blossom Products sold 17,920 DVDs and 4,480 equipment sets. Information on the two products is as follows:

DVDs

Equipment

Sets

Price $8.20 $25.50
Variable cost per unit 4.30 15.20

Total fixed cost is $108,780.

Required:
1. What is the sales mix of DVDs and equipment sets?
2. Compute the break-even quantity of each product.

X

Sales Mix and Break-Even Quantity

1. What is the sales mix of DVDs and equipment sets?

Select4:1

2. Compute the break-even quantity of each product.

The break-even quantity for DVDs is .

The break-even quantity for equipment sets is .
Answer #1

1) Sales mix = 17,920 : 4480 = 4 : 1

2)To calculate Break even quantity we need combined CM Margin.

Particular DVD Equipment
sale price $8.2 $25.50
variable cost 4.3 15.2
Contribution per unit 3.9 $10.3

Combined cm = 14.2..

BREAK EVEN QUANTITY = FIXED COST / Combined CM => $108,780/14.2 = 7660 Units.

DVD BEQ is 7660 * 4/5 = 6128 units

Equipment set BEQ is 7660 * 1/5 = 1532 units.

IFyou need additional data please comment